# Exceptional Points in the Scattering Resonances of a Sphere Dimer

### v1

@@ V Conclusion
In the quasistatic limit, we have established a $\mathscr{P}\mathscr{T}$-symmetric realization of an EPD; consisting of two spheres with complex-conjugate susceptibilities. Beyond this quasistatic limit, retardation breaks $\mathscr{P}\mathscr{T}$-symmetry yet real-frequency EPDs can still be attained by appropriate choices of the dispersion of the materials of the two spheres and geometry. Casting the problem as an equivalent two-port circuit yields analytical criteria for EPDs, whose predictions benchmarked against rigorous multiparticle Mie theory for Drude metals, quantitatively capture the evolution and coalescence of the dimer resonances.
We have characterized the scattering, extinction, and absorption efficiencies across the EPD transition. Close to an EPD, a single-parameter perturbation splits a resonance in the power spectrum into two efficiency peaks whose separation follows the characteristic square-root law seen for the eigenfrequencies. This square root splitting is the EPD-hallmark pointing to enhanced-sensitivity nanophotonic sensors based on minimal building blocks, with tunability via gain/loss control, plasma-frequency engineering, and interparticle spacing.
The framework is readily extensible: higher multipolar orders occurring at larger size parameters enter through additional radiation impedances without altering the synthesis workflow; more complex dispersion can be incorporated via refined material impedances. We anticipate that the same methodology will generalize to complex arrays of scatterers, providing a systematic route to engineer, analyze, and exploit EPDs in realistic photonic platforms.
@@ Appendix C COUPLED-DIPOLE APPROXIMATION
In the limit of small and well-separated particles, and assuming that the real part of their permittivity is either negative or, if positive, not excessively large, the dominant contribution arises from the electric dipole modes, each with impedance $\mathcal{Z}^{h}=\mathcal{Z}^{\mathrm{ed}}$ ($h=1,2,3$). This yields the approximation:
@@
The average current $\langle\mathbf{J}_{\mu}\rangle$ is related to the electric dipole moment $\mathbf{p}_{\mu}$ of particle $\mu$ as
@@
Notably, using the RC approximation of $\mathcal{Z}^{\mathrm{ed}}$ [see Eq. (46)] in Eq. (44) returns the Clausius-Mossotti polarizability with the radiative correction
$\alpha_{\mu}\approx 4\pi a^{3}\left[\frac{\chi_{\mu}}{\chi_{\mu}+3}+j\frac{2}{3}(k_{0}a)^{3}\right].$ (45)
@@ Appendix D ASYMPTOTIC EXPANSIONS
The radiation impedance of the electric dipole mode admits a Laurent series expansion around $k_{0}a=0$. Accordingly, the radiation resistance $\mathcal{R}^{\text{ed}}$ and the modal reactance $\mathcal{X}^{\text{ed}}$ of the electric dipole mode of a sphere can be written as [23]
@@
The mutual resistance $R_{M}^{h}=\Re\{Z_{M}^{h}\}$ and reactance $X_{M}^{h}=\Im\{Z_{M}^{h}\}$ are
@@
$R_{M}^{\overset{y}{z}}$ $=\frac{2}{3}\zeta_{0}\rho^{2}\,\left([(k_{0}d)^{2}-1]\frac{\sin(k_{0}d)}{k_{0}d}+\cos(k_{0}d)\right);$ (48c)
$X_{M}^{\overset{y}{z}}$ $=\frac{2}{3}\zeta_{0}\rho^{2}\,\left([(k_{0}d)^{2}-1]\frac{\cos(k_{0}d)}{k_{0}d}-\sin(k_{0}d)\right).$ (48d)
@@ Appendix E STATE-SPACE EQUATIONS AND EPDs
$\dot{\mathbf{x}}=A\,\mathbf{x}+B\,\mathbf{e},\qquad\mathbf{i}=C\,\mathbf{x}+D\,\mathbf{e},$ (51)
where $\mathbf{x}=\mathbf{x}(t)\in\mathbb{R}^{N}$ is the state vector, $\mathbf{e}(t)=[e_{1},e_{2}]^{\intercal}$ is the input vector, and $\mathbf{i}(t)=[i_{1},i_{2}]^{\intercal}$ is the output vector. The matrix $A\in\mathbb{R}^{N\times N}$ is the system (state) matrix, while $B$, $C$, and $D$ are the input, output, and feedthrough matrices. Since the circuit exhibits no instantaneous dependence of the output on the input, one has $D=0$.
Without any loss of generality we set $e_{2}=0$, as a consequence the the circuit in Fig. 7(a) becomes the one shown in Fig. 7(b) , which is governed by the same state matrix $A$. The corresponding state–space equations read
$\dot{\mathbf{x}}=A\,\mathbf{x}+\mathbf{b}\,e_{1},\qquad i_{1}(t)=\mathbf{c}^{\intercal}\mathbf{x}.$ (52)
where $\mathbf{b}$ and $\mathbf{c}$ are the column vectors associated to the first rows of matrices $B$ and $C$, respectively.
@@
$I_{1}(s)=Y(s)\,\mathcal{E}_{1}(s),\qquad Y(s)=\mathbf{c}^{\intercal}(s\mathbb{I}-A)^{-1}\mathbf{b}.$ (53)
Here, $s=\sigma+j\omega$ denotes the Laplace variable and $\mathbb{I}$ is the identity matrix. The function $Y(s)$ is the transfer function of the system.
@@
so that $Y(s)=Z_{\mathrm{in}}^{-1}(s)$. According to Eq.  (53), if $\lambda_{i}$ is a pole of $Y(s)$ (equivalently, a zero of $Z_{\mathrm{in}}(s)$), then $\lambda_{i}$ is an eigenvalue of $A$.
Moreover, if $Y(s)$ has a pole of order $n_{i}$ at $s=\lambda_{i}$, then the Jordan canonical form of $A$ contains a Jordan block associated with $\lambda_{i}$ of dimension at least $n_{i}$. For a minimal realization, i.e., in the absence of pole-zero cancellations, the dimension of the largest Jordan block associated with $\lambda_{i}$ is exactly $n_{i}$ [27].
As a direct consequence, the presence of higher-order poles $(n_{i}\geq 2)$ in the transfer function $Y(s)=1/Z_{\mathrm{in}}(s)$ implies that the state matrix $A$ is non-diagonalizable. Such spectral degeneracies correspond to exceptional point degeneracies in the dynamical description of the circuit.
@@ Appendix F PARITY-TIME TRANSFORMATION
We now show that the $\mathscr{P}\mathscr{T}$-symmetry implies condition (16). Let $F(\omega):=\det\!\bigl(\boldsymbol{\mathcal{Z}}(\omega)\bigr)$. Under the action of the $\mathscr{P}\mathscr{T}$-operator, the impedance matrix transforms as: $\mathscr{P}\mathscr{T}\{\boldsymbol{\mathcal{Z}}(\omega)\}=-\,\mathbb{P}\,\boldsymbol{\mathcal{Z}}^{\,*}(\omega^{*})\,\mathbb{P}^{\!\intercal}$. Taking determinants and using the $\mathscr{P}\mathscr{T}$-invariance of the system yields $F(\omega;\boldsymbol{\vartheta})=\det\!\left(-\mathbb{P}\,\boldsymbol{\mathcal{Z}}^{\,*}(\omega^{*})\,\mathbb{P}^{\!\intercal}\right)$. Since $\det(\mathbb{P})=1$, we obtain $F(\omega;\boldsymbol{\vartheta})=\det\!\bigl(\boldsymbol{\mathcal{Z}}(\omega^{*})\bigr)^{*}=\bigl[F(\omega^{*};\boldsymbol{\vartheta})\bigr]^{*}$. Thus, $F(\omega^{*};\boldsymbol{\vartheta})=\bigl[F(\omega;\boldsymbol{\vartheta})\bigr]^{*}$, which proves the claim.
@@ Appendix G OPTICAL CROSS-SECTIONS
The time-average powers scattered ($P_{\mathrm{s}}$), absorbed ($P_{\mathrm{a}}$) and extinguished ($P_{\mathrm{e}}$) by the dimer embedded in vacuum are [30]
$P_{\mathrm{s}}=-\frac{V_{a}}{2}\Re\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}^{\mathrm{s}}_{\mu}\rangle,$ (56a)
$P_{\mathrm{a}}=\frac{V_{a}}{2}\Re\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}_{\mu}\rangle,$ (56b)
$P_{\mathrm{e}}=\frac{V_{a}}{2}\Re\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}^{\mathrm{i}}_{\mu}\rangle.$ (56c)
The normalization of these powers by $P_{\mathrm{i}}=\pi a^{2}|\mathbf{E}^{\mathrm{i}}|^{2}/(4\zeta_{0})$, defines the extinction ($Q_{\mathrm{e}}$), scattering ($Q_{\mathrm{s}}$), and absorption ($Q_{\mathrm{a}}$) efficiencies [30]. In the CDA, these three powers can be rewritten as
$P_{\mathrm{s}}\approx\frac{1}{2}\Re\sum_{h=1}^{3}\mathbf{I}^{h\dagger}\begin{bmatrix}\mathcal{Z}^{\mathrm{ed}}&Z_{M}^{h}\\ Z_{M}^{h}&\mathcal{Z}^{\mathrm{ed}}\end{bmatrix}\mathbf{I}^{h},$ (57a)
$P_{\mathrm{a}}\approx\frac{1}{2}\Re\sum_{h=1}^{3}\mathbf{I}^{h\dagger}\begin{bmatrix}Z_{1}^{\mathrm{m}}&0\\ 0&Z_{2}^{\mathrm{m}}\end{bmatrix}\mathbf{I}^{h},$ (57b)
@@ Appendix H PERTURBATION ANALYSIS
An immediate implication of Eq. (58) for the power spectrum is that the resonance frequency, $\omega_{\mathrm{r}\pm}=\Re\{\omega_{\pm}\}$, is highly sensitive to perturbations. In particular, the resonance frequency splitting, $\Delta\omega_{\mathrm{r}}=\omega_{\mathrm{r+}}-\omega_{\mathrm{r-}}$, is given by
$\Delta\omega_{\mathrm{r}}=\begin{cases}-2\sqrt{\delta\vartheta_{i}}\,\Im\{b\},&\delta\vartheta_{i}\geq 0,\\[6.0pt] -2\sqrt{|\delta\vartheta_{i}|}\,\Re\{b\},&\delta\vartheta_{i}<0.\end{cases}$ (59)
@@
$S_{\vartheta_{i}}\{\Delta\omega_{\mathrm{r}}^{\pm}\}:=\left|\frac{\partial\Delta\omega_{\mathrm{r}}}{\partial\vartheta_{i}}\right|_{\vartheta_{i}=\vartheta_{i,\mathrm{ep}}}.$ (60)

### v2

@@ V Conclusion
In the quasistatic limit, we have established a $\mathscr{P}\mathscr{T}$-symmetric realization of an EPD, consisting of two spheres with complex-conjugate susceptibilities. Beyond this quasistatic limit, retardation breaks $\mathscr{P}\mathscr{T}$-symmetry; yet real-frequency EPDs can still be attained by appropriate choices of the dispersion of the materials of the two spheres and geometry. Casting the problem as an equivalent two-port circuit yields analytical criteria for EPDs; the resulting predictions, benchmarked against rigorous multiparticle Mie theory for Drude metals, quantitatively capture the evolution and coalescence of the dimer resonances.
We have characterized the scattering, extinction, and absorption efficiencies across the EPD transition. Close to an EPD, a single-parameter perturbation splits a resonance in the power spectrum into two efficiency peaks whose separation follows the characteristic square-root law seen for the eigenfrequencies. This square-root splitting is a hallmark of EPDs and points to enhanced-sensitivity nanophotonic sensors based on minimal building blocks, with tunability via gain/loss control, plasma-frequency engineering, and interparticle spacing. Achieving the EPD condition exactly remains challenging, particularly at optical frequencies. By contrast, its implementation may be more practical in the microwave regime, where the two Drude permittivities can be implemented by metamaterials, with the active element realized through a tunable negative-resistance load (see, e.g., [36]). The theory can also be applied to anisotropic polarizable particles at microwave frequencies, like dipole antennas connected to an active negative-resistance component.
The framework is readily extensible: higher multipolar orders occurring at larger size parameters enter through additional mode impedances without altering the synthesis workflow; more complex dispersion can be incorporated via refined material impedances. We anticipate that the same methodology will generalize to complex arrays of scatterers, providing a systematic route to engineer, analyze, and exploit EPDs in realistic photonic platforms.
@@ Appendix C COUPLED-DIPOLE APPROXIMATION
In this section, we demonstrate the equivalence of our circuit model to the coupled-dipole approximation model [23, 37, 22, 38, 18].
@@
In the limit of small and well-separated particles, and assuming that the real part of their permittivity is either negative or, if positive, not excessively large, the dominant contribution arises from the electric-dipole modes, each with impedance $\mathcal{Z}^{h}=\mathcal{Z}^{\mathrm{ed}}$ ($h=1,2,3$). This yields the approximation:
@@
The average current $\langle\mathbf{J}_{\mu}\rangle$ is related to the electric-dipole moment $\mathbf{p}_{\mu}$ of particle $\mu$ as
@@
Notably, using the RC approximation of $\mathcal{Z}^{\mathrm{ed}}$ (see Eq. (46), with $\mathcal{X}^{\mathrm{ed}}$ truncated at the order $\omega^{-1}$) in Eq. (44) returns the Clausius-Mossotti polarizability with the radiative correction [23, 22, 18]
$\alpha_{\mu}\approx 4\pi a^{3}\left[\frac{\chi_{\mu}+3}{\chi_{\mu}}+j\frac{2}{3}(k_{0}a)^{3}\right]^{-1}.$ (45)
Accordingly, the proposed circuit model is equivalent to the standard CDA model, but with the polarizability in (44) replacing the Clausius–Mossotti or long-wavelength-approximation polarizability typically adopted in CDA [23, 22, 18]. Nevertheless, Eq. (44) also reveals a key distinction between the two formulations. In the standard CDA, material properties, scattering properties, geometry, and frequency dependence are all incorporated into a single effective polarizability. In contrast, the eigenvalue formulation in Eq. (6) isolates the material contribution from the geometric and radiative terms, which substantially simplifies the derivation of the EPD conditions and adds physical insight.
@@ Appendix D ASYMPTOTIC EXPANSIONS
The impedance of the electric-dipole mode admits a Laurent series expansion around $k_{0}a=0$. Accordingly, the radiation resistance $\mathcal{R}^{\text{ed}}$ and the modal reactance $\mathcal{X}^{\text{ed}}$ of the electric-dipole mode of a sphere can be written as [28]
@@
The mutual resistance $R_{M}^{h}=\operatorname{Re}\{Z_{M}^{h}\}$ and reactance $X_{M}^{h}=\operatorname{Im}\{Z_{M}^{h}\}$ are
@@
$R_{M}^{\overset{y}{z}}$ $=\frac{1}{3}\zeta_{0}\rho^{2}\,\left([(k_{0}d)^{2}-1]\frac{\sin(k_{0}d)}{k_{0}d}+\cos(k_{0}d)\right);$ (48c)
$X_{M}^{\overset{y}{z}}$ $=\frac{1}{3}\zeta_{0}\rho^{2}\,\left([(k_{0}d)^{2}-1]\frac{\cos(k_{0}d)}{k_{0}d}-\sin(k_{0}d)\right).$ (48d)
@@ Appendix E STATE-SPACE EQUATIONS AND EPDs
$\dot{\mathbf{x}}=\boldsymbol{A}\mathbf{x}+\boldsymbol{B}\mathbf{e},\qquad\mathbf{i}=\boldsymbol{C}\mathbf{x}+\boldsymbol{D}\mathbf{e},$ (51)
where $\mathbf{x}=\mathbf{x}(t)\in\mathbb{R}^{N}$ is the state vector, $\mathbf{e}(t)=[e_{1},e_{2}]^{\intercal}$ is the input vector, and $\mathbf{i}(t)=[i_{1},i_{2}]^{\intercal}$ is the output vector. The matrix $\boldsymbol{A}\in\mathbb{R}^{N\times N}$ is the system (state) matrix, while $\boldsymbol{B}$, $\boldsymbol{C}$, and $\boldsymbol{D}$ are the input, output, and feedthrough matrices. Since the circuit exhibits no instantaneous dependence of the output on the input, one has $\boldsymbol{D}=\boldsymbol{0}$.
Without loss of generality, we set $e_{2}=0$, as a consequence the circuit in Fig. 7(a) becomes the one shown in Fig. 7(b), which is governed by the same state matrix $\boldsymbol{A}$. The corresponding state–space equations read
$\dot{\mathbf{x}}=\boldsymbol{A}\mathbf{x}+\mathbf{b}\,e_{1},\qquad i_{1}(t)=\mathbf{c}^{\intercal}\mathbf{x}.$ (52)
where $\mathbf{b}$ is the column vector corresponding to the first column of $\boldsymbol{B}$, and $\mathbf{c}$ is the column vector whose transpose $\mathbf{c}^{\intercal}$ corresponds to the first row of $\boldsymbol{C}$.
@@
$I_{1}(s)=Y(s)\,\mathcal{E}_{1}(s),\qquad Y(s)=\mathbf{c}^{\intercal}(s\mathbb{I}-\boldsymbol{A})^{-1}\mathbf{b}.$ (53)
Here, $I_{1}$ and $\mathcal{E}_{1}$ denote the Laplace transforms of $i_{1}$ and $e_{1}$, respectively; $s=\sigma+j\omega$ denotes the Laplace variable, and $\mathbb{I}$ is the identity matrix. The function $Y(s)$ is the transfer function of the system.
@@
so that $Y(s)=Z_{\mathrm{in}}^{-1}(s)$. According to Eq.  (53), if $\lambda_{i}$ is a pole of $Y(s)$ (equivalently, a zero of $Z_{\mathrm{in}}(s)$), then $\lambda_{i}$ is an eigenvalue of $\boldsymbol{A}$.
Moreover, if $Y(s)$ has a pole of order $n_{i}$ at $s=\lambda_{i}$, then the Jordan canonical form of $\boldsymbol{A}$ contains a Jordan block associated with $\lambda_{i}$ of dimension at least $n_{i}$. For a minimal realization, i.e., in the absence of pole-zero cancellations, the dimension of the largest Jordan block associated with $\lambda_{i}$ is exactly $n_{i}$ [33].
As a direct consequence, the presence of higher-order poles $(n_{i}\geq 2)$ in the transfer function $Y(s)=1/Z_{\mathrm{in}}(s)$ implies that the state matrix $\boldsymbol{A}$ is non-diagonalizable. Such spectral degeneracies correspond to exceptional point degeneracies in the dynamical description of the circuit.
@@ Appendix F PARITY-TIME TRANSFORMATION
We now show that the $\mathscr{P}\mathscr{T}$-symmetry implies condition (16). Let $F(\omega):=\det\!\bigl(\boldsymbol{\mathcal{Z}}(\omega)\bigr)$. Under the action of the $\mathscr{P}\mathscr{T}$-operator, the impedance matrix transforms as: $\mathscr{P}\mathscr{T}\{\boldsymbol{\mathcal{Z}}(\omega)\}=-\,\mathbb{P}\,\boldsymbol{\mathcal{Z}}^{*}(\omega^{*})\,\mathbb{P}^{\intercal}$. Taking determinants and using the $\mathscr{P}\mathscr{T}$-invariance of the system yields $F(\omega;\boldsymbol{\vartheta})=\det\!\left(-\mathbb{P}\,\boldsymbol{\mathcal{Z}}^{*}(\omega^{*})\,\mathbb{P}^{\intercal}\right)$. Since $\det(\mathbb{P})=-1$, we obtain $F(\omega;\boldsymbol{\vartheta})=\det\!\bigl(\boldsymbol{\mathcal{Z}}(\omega^{*})\bigr)^{*}=\bigl[F(\omega^{*};\boldsymbol{\vartheta})\bigr]^{*}$. Thus, $F(\omega^{*};\boldsymbol{\vartheta})=\bigl[F(\omega;\boldsymbol{\vartheta})\bigr]^{*}$, which proves the claim.
@@ Appendix G OPTICAL CROSS-SECTIONS
The time-averaged powers scattered ($P_{\mathrm{s}}$), absorbed ($P_{\mathrm{a}}$) and extinguished ($P_{\mathrm{e}}$) by the dimer embedded in vacuum are [39]
$P_{\mathrm{s}}=-\frac{V_{a}}{2}\operatorname{Re}\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}^{\mathrm{s}}_{\mu}\rangle,$ (56a)
$P_{\mathrm{a}}=\frac{V_{a}}{2}\operatorname{Re}\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}_{\mu}\rangle,$ (56b)
$P_{\mathrm{e}}=\frac{V_{a}}{2}\operatorname{Re}\sum_{\mu=1}^{2}\langle\mathbf{J}_{\mu}|\mathbf{E}^{\mathrm{i}}_{\mu}\rangle.$ (56c)
The normalization of these powers by $P_{\mathrm{i}}=\pi a^{2}|\mathbf{E}^{\mathrm{i}}|^{2}/\zeta_{0}$, defines the extinction ($Q_{\mathrm{e}}$), scattering ($Q_{\mathrm{s}}$), and absorption ($Q_{\mathrm{a}}$) efficiencies [39]. In the CDA, these three powers can be rewritten as
$P_{\mathrm{s}}\approx\frac{1}{2}\operatorname{Re}\sum_{h=1}^{3}\mathbf{I}^{h\dagger}\begin{bmatrix}\mathcal{Z}^{\mathrm{ed}}&Z_{M}^{h}\\ Z_{M}^{h}&\mathcal{Z}^{\mathrm{ed}}\end{bmatrix}\mathbf{I}^{h},$ (57a)
$P_{\mathrm{a}}\approx\frac{1}{2}\operatorname{Re}\sum_{h=1}^{3}\mathbf{I}^{h\dagger}\begin{bmatrix}Z_{1}^{\mathrm{m}}&0\\ 0&Z_{2}^{\mathrm{m}}\end{bmatrix}\mathbf{I}^{h},$ (57b)
@@ Appendix H PERTURBATION ANALYSIS
An immediate implication of Eq. (58) for the power spectrum is that the resonance frequency, $\omega_{\mathrm{r}\pm}=\operatorname{Re}\{\omega_{\pm}\}$, is highly sensitive to perturbations. In particular, the resonance frequency splitting, $\Delta\omega_{\mathrm{r}}=\omega_{\mathrm{r+}}-\omega_{\mathrm{r-}}$, is given by
$\Delta\omega_{\mathrm{r}}=\begin{cases}-2\sqrt{\delta\vartheta_{i}}\,\operatorname{Im}\{b\},&\delta\vartheta_{i}\geq 0,\\[6.0pt] -2\sqrt{|\delta\vartheta_{i}|}\,\operatorname{Re}\{b\},&\delta\vartheta_{i}<0.\end{cases}$ (59)
@@
$S_{\vartheta_{i}}\{\Delta\omega_{\mathrm{r}}\}:=\left|\frac{\partial\Delta\omega_{\mathrm{r}}}{\partial\vartheta_{i}}\right|.$ (60)
@@
Appendix I TRANSVERSE POLARIZATION
In Sec. IV, results are presented for the longitudinal polarization of the dipolar mode ($h=x$). In this appendix, we discuss the transverse polarization case ($h=y,z$).
The primary difference arises from the coupling mechanism between the dipoles. Within the circuit-based formalism, this modifies only the mutual impedance $Z_{M}$. This difference becomes particularly transparent in the quasistatic limit. In this regime, Eq. (22) becomes
$\left(\frac{\omega}{\omega_{0}}\right)^{4}+\left(\hat{\gamma}^{2}-2\right)\left(\frac{\omega}{\omega_{0}}\right)^{2}+1-\rho^{6}=0,$ (61)
The corresponding eigenfrequencies are
$\omega_{\pm}(\hat{\gamma},\rho)=\omega_{0}\sqrt{1-\frac{\hat{\gamma}^{2}}{2}\pm\frac{1}{2}\sqrt{\hat{\gamma}^{4}-4\hat{\gamma}^{2}+4\rho^{6}}}.$ (62)
An EPD occurs when the inner discriminant vanishes:
$\hat{\gamma}^{4}-4\hat{\gamma}^{2}+4\rho^{6}=0,$ (63)
which is quadratic in $\hat{\gamma}^{2}$. The corresponding discriminant is $\Delta=16(1-\rho^{6})$, which remains strictly positive under the non-overlapping condition $d>2a$ (i.e., $\rho<1/2$). Therefore, as in the longitudinal case, two real solutions for $\hat{\gamma}^{2}$ exist throughout the physically admissible parameter range.
Compared to (25), Eq. (63) differs only in the coefficient of the $\rho^{6}$ term.